OpenStudy (anonymous):

23. Show all work to simplify 2 over x plus 2 over x – 1 minus 2 over x – 2. Use complete sentences to explain how to simplify this expression. Remember to list all restrictions.

OpenStudy (anonymous):

\[\frac{ 2 }{ x }+\frac{ 2 }{ x-1 }-\frac{ 2 }{ x-2 }\]

OpenStudy (anonymous):

ill leave the sentences for u to make up lol but i can show u how to simplify this expression ;)

OpenStudy (anonymous):

so wat would you say id the common denominator here?

OpenStudy (anonymous):

yeah i have that set up and the common denominator which is x-2

OpenStudy (anonymous):

wait never mind lol i miss looked at that part

OpenStudy (anonymous):

ok lol want me to tell u or u can do it now?

OpenStudy (anonymous):

I think its x(x-1)(x-2)

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

:) okay lol i am happy i got that i am good at algebra just writing these out is where i get stuck

OpenStudy (anonymous):

makes sense lol :)

OpenStudy (anonymous):

but i got the answer which is